Prev Next

SW&IT Meeting Minutes 2018-06-04

Attendees: Andrew, David, Doug, Igor, Joe, Kellee, Matthieu, Petr, Stephen, Xianyu, Yang


News from management

  • xxx

Previous Action Items

  • None

Open topics

Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.

  • LUCI Sun [MB]
  • revision control workflow [IS]
  • coding standards, indent tools (discussion) [MB]
  • ops-support oriented `single` wiki page (do we need it?) [MB]
  • SSD plan w. Leroy and Chris [MB]
  • git migration, next? [MB]
  • git local cron clone [MB]
  • VMware license
  • AC/YZ support July 2018
  • Team calendar / UAccess, ways to keep it simple

This week in github

Team reports


  • MPIA TRAC conversion to LBTO wiki (85% complete)
  • Updated various wiki pages: data dictionary list, WFS camera control, etc.
This week:
  • Handover zoom with Wolfgang, Jose, and Matthieu
  • Finish wiki conversion
  • First final draft of handover document


  • Working on IT email problem; finalizing document today
  • Miscellaneous tasks like setting up laptop for Kara, working with Dell to transfer ARGOS computers to us in the US + service contract, Obs room UPS, etc.
This week:
  • With approval, begin modifications to IT distribution lists
  • Research VMWare upgrades
  • Backup testing


Last 2 Weeks
  • HR activities
  • Transition reading materials prep (Yang, Petr, Matthieu)
  • Some initial email dialog on pointing & collimation with Petr & Yang.
  • Laptop Procurement & transition prep
  • Monitoring of Pointing related ITs 7269 (closed), + 7276
  • Evaluation of IE/CA changes for facility instruments over last month+ & implications discussion with Mark W.
This week:
  • Laptop transition continues
  • Pointing & Collimation transition discussions continued (should Petr and Yang be available).
  • Short week this week - Wed is Vac, so only Mon & Tues are available.


OT and PIT
  • PIT spreadsheets for italian partner
  • Starting AFC OT planning/design with Michelle and Dave T.. Doing some scouting on the code to see if some of the ideas put forward are easily implementable withing the current OT architecture.

  • LUCI scheduler new interface. I've been having a lot of second guessing on the design, still not happy but I have to call it a day. Wish it were more MVC oriented but not quite there, maybe will try that using a Full Stack aproach. (by later on, for now will roll with what we have)
  • Create SQL to populate AFC table with new set of AFC masks.
  • Improve on the shell script, hopefully will make part of the LUCI removal procedure. In fact ideally we should create software so that the engineer in charge of doing the LUCI removal can do that without software team assistance.


Last 2 weeks
  • Six day vacation and a holiday.
  • Discuss migration of OT trunk/ocs to Github with MB. This has been deprioritized for now, so I will continue using a dev branch on the existing Subversion archive.
  • Discuss Planner with ME+IS. (DONE, notes follow)
  • Steps should be labeled INST - [OB#] - NNN (DONE)
  • Each Sync Offset should be toggle selectable on the Step Plotter (DONE)
  • Sync Offset steps should check copoint from offset position not preset (DONE)
  • OBs should plan directly from OT, no more drag + drop between ready and planned lists. (Partial, I added a menu item that does this but I think we want drag + drop)
  • Step plotter should display mirror targets as well as mount position
  • Steps should be computed including mount offset (DONE)
  • Change to mirror targets should update steps
  • Tooltip should contain times to complete sequence items in addition to their values, and the sum of the time used.
  • Planner should be reentrant between science plans
This week:
  • Finish any uncompleted items


Last two weeks:
  • OAC HK upgrade
    • Completed the changes to OAC HK upgrade for providing HK data even if UMAC is unavailable.
    • Got new OAC version built in 32bit environment in Tucson.
    • Started the GCS updates for the OAC HK upgrade.
    • Revised OAC to no longer lump UMAC data with onewire and housekeeping data - new RPC client, etc.
  • Created a 2.15 LBC software release with a fix for IT5715. Sci Ops didn't get to it, so it has not been released.
  • Reviewed updates to DDS by MB.
  • Added the MODS handover document to the VCAN (I603s00201). Pinged Rick and Jerry about next steps.
  • Modified the agw4-cam version of azcamserver in the Tucson lab to NOT overwrite the exposure each time for Jonathan's testing.
  • Modified GCS to use threaded version of asctime (IT7045).

Next week:
  • Prepare OVMS racktangle kernel update - recompile with newer files.
  • Continue GCS updates for new OAC HK.
  • Support engineering team tests of OAC HK upgrade.


  • management OC, WPC
  • Engineering manager interviews
  • S/W handovers (MODS, ARGOS, LUCI) follow up
  • USB3 extenders follow up
  • LUCI Sun follow up
  • PEs (continued)
  • iLocater, SHARK-NIR/VIS monthly
  • obsdev meeting
  • LBC Ethernet controller, Fernando at Startech week after SPIE
  • PBP meeting
  • Ketiv/Vault iteration

  • return PEs
  • handovers follow up
  • ETC with Mark and Al
  • responsibilities and assignments


  • vacations, worked remotely on copointing
  • working on copointing change, nearing solution (quite few obstacles on the path to newer, better, simpler code removed; works at the moment except for checks after absorbing offsets, need to investigate what is happening)
  • TCSRelease wiki page

Next week:
  • collimation study


  • Docker build images, auto-provisioning for build (not complete)
  • Prep for Network tech talk.
  • AO system kickstart attempt (automated install of AO systems)
  • Microwave data storage and presentation (about an hour of scripting, and an hour of playing with Grafana/MySQL).
  • help with AGW4 in the lab.
  • Google drive shared folders (setup)
  • Account cleanup in Google (what to do with employees who have left?)
  • Changes to conference room reservations.
  • Physical inventory concerns.
  • Account setup for Kara.
  • Quote and purchase of two ASAs for the mountain.

  • Continue with Docker, cisrv01, docker registry (which seems a little evil)
  • Github hooks into cisrv01
  • AGW1 assist (likely)
  • Custom widgets in Grafana?
  • Containerized apps.


Late 2 weeks:
  • Reviewed and sent comments to MG for PBP memory map V1 and V2
  • Meeting with MG about PBP
  • Discussed/showed the PBP SW changes with MB, JC and LF
  • Configured the spare Adsec box for power backplane, recorded by stephen for kickstart
  • Worked on the SOUL box requirement and the order
  • Tested and installed the shell safety in AdamhouseKeeper
  • Tested and installed the changes in Adsec Arbitrator
  • Discussions with Guido and Doug M. for the AdamhouseKeeper and Adsec Arbitrator changes
  • Tested and installed the AdamhouseKeeper and Adsec Arbitrator changes
  • Update IT7260 and installed the small change
  • Collected the current housekeeper and fastdiagnostic actions
This week:
  • Working on the PBP


  • Temporary solution submitted to IT issue #7260 (Msgd connection problem).
  • Working on IT issue #7261 (not exactly sure what the problem is, or in fact it could be due to general poor design, which will be nearly impossible to fix, or take very long time).
  • Working on wiki documentation on AO software subsystem and logging outputs (needs collaborative editing from people familiar with AO software system).
  • Doing some more AO simulator related study and testing (include logging).
* Study pointing model related material. * Study pointing model related material.
Topic revision: r20 - 10 Dec 2018, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback